Find the standard deviation of 1, 2, 3, 3, 5, 10. (a) 2.66 (b) 3.17 (c)...

  1. Find the standard deviation of 1, 2, 3, 3, 5, 10.

    1. (a) 2.66

    2. (b) 3.17

    3. (c) 3

    4. (d) 2.94
Answer #1

d.2.94

first let us know the mean of given numbers

=>(1+2+3+3+5+10) /6

=>4.

standard deviation = \sqrt{}[sum of ( observation - mean)^2] / number of observations

standard deviation = \sqrt{}[(1-4)^2 + (2-4)^2+(3-4)^2+(3-4)^2+(5-4)^2+(10-4)^2] / 6

=>\sqrt{}(9+4+1+1+1+36)/6

=>\sqrt{}52/6

=>2.94.